OpenStudy (anonymous):

find the value of x that makes quadrilateral ABCD a rectangle , if BE=3x+1, and ED=5x-3 A) 3 B) 2 C) 5 D)7

OpenStudy (nurali):

3x+1=5x-3 5x-3x=3+1 2x=4 x=2